14,20,26 10,24,38 44,28,12 22,?,14
Answer by math_helper(2461)

The are 4 groups of 3 numbers each:

14,20,26 <— notice 20 = (14+26)/2
10,24,38 <— 24 = (38+10)/2
etc.
following this pattern ? = (22+14)/2 = _____________ (you turn the crank)